OpenStudy (anonymous):

find the sum of the infinite series" 12-1+1/12-1/144+...

OpenStudy (anonymous):

\[\frac{a}{1-r}\] for this one

OpenStudy (perl):

hint, you're series involves powers of 12

OpenStudy (anonymous):

evidently \(a=12\)

OpenStudy (anonymous):

and \(r=?\)

OpenStudy (anonymous):

is it clear what \(r\) is in this example?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

r is the ratio of one term to the previous term, not the second term

OpenStudy (anonymous):

take any term and divide it by the previous one it should really be obvious from your eyeballs, and not require a computation what do you multiply 12 by to get -1?

OpenStudy (anonymous):

negative 12 @satellite73

OpenStudy (anonymous):

close but you are doing it backwards

OpenStudy (anonymous):

\[12\times r=-1\]

OpenStudy (anonymous):

\[r=?\]

OpenStudy (anonymous):

is that even possible? @satellite73

OpenStudy (anonymous):

i sure hope so because you are doing it an infinite number of times

OpenStudy (anonymous):

all you have to do is write it \[r=-\frac{1}{12}\]

OpenStudy (anonymous):

12/(1 - (-1/12)) = 12/(13/12) = 144/13 = 11.076?? @satellite73

OpenStudy (anonymous):

i would leave it at \(\frac{144}{13}\) looks a lot prettier good job though